# Exact Work Distribution and Jarzynski’s Equality of a Relativistic Particle in an Expanding Piston

### v1

@@ Appendix B Details of the Integration
$T_{n}=\frac{\alpha_{p}^{-(n-1)}-1-\alpha_{p}+\alpha_{p}^{n}}{1-\alpha_{p}}+\frac{\alpha_{p}^{-(n-1)}+\alpha_{p}^{n}}{1+\alpha_{p}}\cdot\tau$ (50)
@@
$X_{n}=\frac{\alpha_{p}^{-(n-1)}-\alpha_{p}^{n}}{1-\alpha_{p}}+\frac{\alpha_{p}^{-(n-1)}-\alpha_{p}^{n}}{1+\alpha_{p}}\cdot\tau$ (51)
are the slopes and the intercepts of the lines respectively. The separation of the domain of the integration is depicted in Fig. 4.
Figure 4: Dividing the domain of integration into parts and meaning of the overlap factor, where the length of the red dashed line is the value of $\varphi_{n}$ at the same $v$
@@
What is left involves integrating a Dirac $\delta$ function, we have
@@
Summing up all the pieces at hand we have the result Eq. (27).

### v2

@@ Appendix D Details of the Integration
$T_{n}=\frac{\alpha_{p}^{-(n-1)}-1-\alpha_{p}+\alpha_{p}^{n}}{1-\alpha_{p}}+\frac{\alpha_{p}^{-(n-1)}+\alpha_{p}^{n}}{1+\alpha_{p}}\tau$ (54)
@@
$X_{n}=\frac{\alpha_{p}^{-(n-1)}-\alpha_{p}^{n}}{1-\alpha_{p}}+\frac{\alpha_{p}^{-(n-1)}-\alpha_{p}^{n}}{1+\alpha_{p}}\tau$ (55)
are the slopes and the intercepts of the lines, respectively. The separation of the domain of the integration is depicted in Fig. 7.
Figure 7: Division the domain of integration into parts. The overlap factor can be explained by the length of the red dashed line being the value of $\varphi_{n}$ at the same $v$.
@@
What is left involves integrating a Dirac $\delta$ function. We have
@@
Summing up all the pieces at hand we have the result (18).